OpenStudy (anonymous):

Can someone explain this to me? Solve the equation by completing the square: 2x2 - 20x = 8

OpenStudy (anonymous):

add 8 both sides and use quadratic formula

OpenStudy (anonymous):

and you get 2x^2-20x-8=0 ??? or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then what?

OpenStudy (anonymous):

2x2 - 20x - 8 = 0 2 ( x2 -10x - 4 ) = 0 2 ( (x-5)^2 - 4 - 5 = 0 2 ( (x-5)^2 - 9 ) = 0 2(x-5)^2- 4.5 = 0

OpenStudy (anonymous):

you can use the quadratic formula

OpenStudy (anonymous):

or try to factor first

OpenStudy (anonymous):

i factor -8?

OpenStudy (anonymous):

did you understand what i did to complete the square ?

OpenStudy (anonymous):

not really >.< this really confuses me.

OpenStudy (anonymous):

im terrible at math

OpenStudy (anonymous):

http://www.webgraphing.com/quadraticequation_completingthesquare.jsp here is your problem

OpenStudy (anonymous):

ok you have 2x^2 - 20x -8 = 0 you cant complete the square when there is a 2 in front of the x^2 so: you remove the 2 outside of the eq. = 2 [ (x^2 -10 x -4 ) = 0 Notice i divide the 20x by 2 and 8 by 2 also... then : 2 [ (x-5)^2 - 4 -5 ) ] Here i divide the 10 by 2 so that I obtain it in a form ( a + b )^2 and I subtract it from the -4 then you will hve 2 [ ( x-5)^2 -9 ] = 0 then you multiply the whole by 2 : 2(x-5)^2- 9/2 = 0 9/2 = 4.5 :) hope I help...

OpenStudy (anonymous):

Thank you, guys :D I appreciate it! (: